# Experimental Study of Bremsstrahlung Gamma Ray Emission and Short-Range Correlations in 124Sn+124Sn Collisions at 25 MeV/u

### v1

@@ IV.1 Model Description
For comparison, a momentum-independent soft Bertsch-Kruse-Das Gupta (SBKD) potential is also used in the IBUU simulations [78],
@@
The bremsstrahlung reaction channel is incorporated into the collision processes within the IBUU model, which employs the Bertsch criterion [87] and Pauli exclusion principle to determine whether a NN collision occurs. Bremsstrahlung photons are generated through two channels, the neutron-proton collisions ($np\gamma$) and the proton-proton collisions ($pp\gamma$). However, the $np\gamma$ channel dominates as the dipole moment in a charge-symmetric $pp$ system vanishes. Consequently, the lowest-order multipole radiation in the $pp$ collisions is of quadrupole type, whose intensity is much lower compared to the dipole radiation in the $np$ collisions [88, 89]. Hence we consider only the $np\gamma$ process in our calculations.
The bremsstrahlung photons emitted in the $np$ collisions are produced in the processes of deceleration and acceleration of the protons interacting with the neutrons. In the intermediate energy range, we adopt the neutral scalar $\sigma$ meson exchange model [88] in the present work, in which the photons are emitted via the external processes, illustrated by the Feynman diagrams in Fig.19 (a). At higher energies, one will have to take the charged meson exchange processes into consideration. The photons can also be emitted from the internal mesons, to which the electromagnetic field also couples [88, 90], and the Feynman diagram is given by Fig.19 (b).
@@
Due to the low probability of producing bremsstrahlung photons in the reaction, the impact of bremsstrahlung on nucleon kinematics is also negligible, and hence, the perturbation method is applied to calculate the probability of photon production in each $np$ collision. Based on the $\sigma$ meson exchange model mentioned above, a well fitted expression for the probability of elementary double differential photon production can be applied in IBUU simulations [91],
@@
where $E_{\gamma}$ represents the energy of produced photons, and $E_{\rm max}$ represents the total available energy in the center of mass frame of the colliding neutron-proton pair. The coefficient $\alpha=0.7319-0.5898\beta$, where $\beta$ represents the velocity of nucleon. The total photon production probability per event is the sum of the probabilities of all $np$ collisions producing photons throughout the entire process.
IV.2 Parameter influence test
@@
In the IBUU simulations, one has to choose several important inputs, which may influence the results of final observables. Here we systematically check the sensitivity of bremsstrahlung photon production on the choices of parameters, namely the impact parameter, the nuclear mean-field potential, and the symmetry energy parameter.
Firstly, we show in Fig. 20 the variations of bremsstrahlung photon production by using different impact parameters. We adopt $b=0\sim 5\,\rm{fm}$ in our IBUU simulations. To check the sensitivity of photon production on $b$ parameter, several other values are used, namely, $b=0\sim 9\,\rm{fm}$ and $b=0,5,9\,\rm{fm}$. Their corresponding spectra are normalized to the $b=0\sim 5\,\rm{fm}$ one. This is because we focus on the spectral shape similarities rather than absolute normalization. Although the absolute values of bremsstrahlung photon production spectra vary with different choices of impact parameters, but the spectral shape remain almost the same to the impact parameters, as seen in Fig. 20.
Figure 20: The double differential photon production calculated by varying impact parameters. The impact parameters are taken to be $b=0\sim 5\,\rm{fm}$, $b=0\sim 9\,\rm{fm}$ and fixed values $b=0,5,9\,\rm{fm}$. The results using $b=0\sim 9,0,5,9\,\rm{fm}$ are normalized to the $b=0\sim 5\,\rm{fm}$ case.
In Fig. 21, the variations of bremsstrahlung photon production by using different nuclear mean-field potentials are shown. We adopt the MDI potential with the symmetry energy parameter x=1x=-1 [66]. The momentum-independent SBKD potential is used to check the sensitivity of mean-field potentials, and the parameters used in SBKD potential are listed in Table3. The $\sigma$ parameter represents the softness or stiffness of the potential with the variation of the nuclear density [78]. For the possible influence of symmetry energy parameters in MDI, we adopt three different parameters x=1x=-1, $x=0$ and $x=1$ in simulations (see Fig. 22). The tests have shown that within the energy region investigated in this work, our conclusions are not affected by the choice of nuclear potentials or the symmetry energy parameters. This is probably because the bremsstrahlung photons interact with nucleons only electromagnetically, and are not sensitive to the isoscalar and isovector parts of nuclear interactions.
@@
Figure 21: The double differential photon production calculated by varying the choice of nuclear mean-field potentials. We compare the results from the MDI potential with those from the momentum-independent SBKD potential. The results from the SBKD potential are normalized to the MDI results.
@@
Table 3: The parameters $\sigma$, $A\,{\rm{(MeV)}}$ and $B\,{\rm{(MeV)}}$ used in the momentum-independent SBKD potential.
$\sigma$ $A\,{\rm{(MeV)}}$ $B\,{\rm{(MeV)}}$
7/6 -356 303
4/3 -218 164
2 -124 70.5
2.77 -103.22 49.56
Figure 22: The double differential photon production calculated by varying symmetry energy parameters. Three different $x$ parameters are employed in the simulations, namely x=1x=-1, $x=0$ and $x=1$. The results using $x=0$ and $x=1$ are normalized to the x=1x=-1 case.
IV.3 Reference Frame Transformation
The mass of each nulceon is $m_{\rm u}=931.5~\rm MeV$ and the kinetic energy of each nulceon is $E_{\rm u}=25~\rm MeV$. The momentum of each nulceon in the beam is
$p_{\rm u}=\sqrt{(E_{\rm u}+m_{\rm u})^{2}-m_{\rm u}^{2}},$ (18)
the rapidity of the beam can be obtained by $p_{u}$ as the following,
$sinhy_{b}=\frac{p_{\rm u}}{m_{\rm u}}.$ (19)
In the system of 124Sn+124Sn, the mass of the nuclei in the beam $M_{\rm b}=124$ and the mass of the target nuclei $M_{\rm t}=124$ can be used for the calculation of the rapidity in the center-of-mass frame system,
$y_{\rm cm}=\frac{1}{2}y_{\rm b}+\frac{1}{2}\ln[\frac{M_{\rm b}\exp(y_{\rm b})+M_{\rm t}}{M_{\rm b}+\exp(y_{\rm b})M_{\rm t}}].$ (20)
And then, we can obtain the velocity in the center-of-mass frame frame,
$\beta_{\rm cm}=\frac{\exp(2y_{\rm cm})-1}{\exp(2y_{\rm cm})+1}.$ (21)
The $\gamma$ detector array is placed at $\theta_{\rm b}=110^{\circ}$ with respect to the beam direction, and the transformation factor can be finally obtained as the following,
$L(\theta_{\rm b})=\frac{1-\beta_{\rm cm}\cos\theta_{\rm b}}{\sqrt{1-\beta_{\rm cm}^{2}}}=1.04631.$ (22)
The photon Doppler effect causes softening of the energy spectrum. During the conversion of the spectrum from the laboratory frame to the center-of-mass frame, the energy must be multiplied by the factor $L$, while the counts are divided by the same factor. The experimentally detected $\gamma$-ray spectrum can be transformed into the center-of-mass frame using this method.
Since the detected $\gamma$-ray spectrum in the experiment is influenced by the detector filter, theoretical curves must also account for the detector filter when being compared with experimental data. Using Geant4 simulations, we modeled the CsI array’s response to $\gamma$-rays of varying energies and obtained the detector filter matrix in the laboratory frame, as depicted in Fig. 16. To incorporate this effect, the theoretical curves were first transformed into the laboratory frame, passed through the detector filter matrix, and then converted back to the center-of-mass frame. This procedure enables a direct comparison between the experimental data and the IBUU theoretical predictions in subsequent analyses.
V Results and Discussions
V.1 Analysis of Systematic Uncertainty
@@
The systematic uncertainties primarily stem from three sources: (1) variations in the spectra obtained from radioactive source calibrations on different dates during the experiment, (2) the choice of normalization energy range for high-energy cosmic rays during background subtraction, and (3) the linear assumption of the detector response at higher $\gamma$ energy levels. Analysis shows that the choice of normalization energy ranges exhibit no significant variations, indicating that the effect of cosmic ray normalization range on the $\gamma$-ray spectrum is negligible.
@@
The experimental data included three sets of linear calibration obtained from radioactive sources on different dates. Among these, the $\gamma$ energy spectrum calibrated using the dataset with the highest statistics was designated as the central spectrum. In addition to the linear calibration, we introduced two alternative detector response correction functions (DRCFs) characterizing the differences between the linear calibration and the actual quadratic calibration obtained on SLEGS test experiment [72]. Then a total of nine distinct $\gamma$ energy spectra can be derived. Excluding the central spectrum, we calculated the standard deviation of the remaining eight spectra for each bin as the systematic uncertainty. During the evaluation of the likelihood functions and $\chi^{2}$ (see next), the values derived from the central spectrum were used as the central points. The likelihood function and $\chi^{2}$ values computed from the remaining spectra were utilized to determine the associated errors (standard deviations). This methodology enabled us to obtain likelihood and $\chi^{2}$ functions with systematic uncertainties at different $R_{\text{HMT}}$ values.
V.2 Experimental $\gamma$ spectrum in slow coincidence
@@
As shown by the primary analysis, the method introduced in the previous experiment [66] is applied. All the $\gamma$ events in the slow coincidence window are counted. Namely a valid timing of the reconstructed $\gamma$ is found within the TDC range. The background is obtained by the beam-off measurement. The experimental data calibration and background subtraction were performed using the background data and radioactive source calibration data on March 8, 2024. Calibration files were generated based on this data and subsequently applied to all experimental and background data for calibration and reconstruction. This process yielded the total beam-on spectrum and the beam-of background spectrum, shown as the black and red lines in Fig. 23(a), respectively. The red background spectrum was scaled by a normalization factor $r_{\rm n}$. Specifically, the total counts above $110\rm MeV$ in both the total spectrum ($Y_{\rm on}$) and the background spectrum ($Y_{\rm off}$) were calculated, and the yield ratio is defined as $r_{\rm n}=\frac{Y_{\rm on}(E_{1})}{Y_{\rm off}(E_{1})}$, where the subscript ‘on’ and ‘off’ represent the beam-on and beam-off spectrum, respectively. $E_{1}=110$ MeV is the low energy border taken to do the normalization, the upper border is set to 200 MeV. Then the ratio $r_{\rm n}$ is applied to the entire background spectrum, resulting in the red histogram in Fig. 23(a). By subtracting the scaled background spectrum from the total energy spectrum, the background-subtracted $\gamma$ spectrum was obtained, as shown in Fig. 23(b), where we take 01000-100 MeV as the energy range of total $\gamma$-rays.
@@
[…]
Figure 24: (Color Online) Comparison of reconstructed energy spectra. (a) Energy spectra reconstructed from the central four units, edge six units, and all units combined. (b) The same spectra after normalization in the $35\sim 100$ MeV energy range.
Fig. 25 illustrates the rebinned experimental $\gamma$ spectrum in the c.m. frame. The black dots represent the central spectrum, with statistical uncertainties indicated by error bars, and the gray shaded regions denote the systematic uncertainties at each energy point. Several key theoretical curves, processed with the detector filter, are presented for comparison. Unlike in our previous measurement [66, 65], in this experiment, the $\gamma$ hodoscope was operated in active triggering mode, preventing the determination of the number of $np$ collisions in heavy-ion reactions as a normalization factor for the $\gamma$ spectrum. Consequently, the comparison between experimental data and theoretical curves predicted by the IBUU model focused on spectral shape similarities rather than absolute normalization. To align the range of the theoretical curves with the experimental spectrum, the theoretical values were uniformly scaled by a factor of $2.5\times 10^{9}$, which had no impact on determining the $R_{\rm HMT}$, as it did not affect the comparison of shape differences between the experimental and model-predicted spectra.
Figure 25: (Color Online) Comparison of the rebinned experimental $\gamma$ spectrum (black dots) with statistical uncertainties (error bars) and systematic uncertainties (gray shaded areas) in the c.m. frame. Several key theoretical curves, processed with the detector filter, are overlaid for comparison.
Energies below $30~\rm MeV$ are primarily influenced by collective resonance and statistical emissions as well as the random coincidence. Therefore, we selected an energy range of $35<E_{\gamma}^{\rm cm}<100~\rm MeV$ as the central analysis interval and calculated the likelihood function between the experimental $\gamma$ spectrum at each point within this range and various IBUU theoretical curves.
@@
where $i$ represents for the sum of experimental points within a certain statistical analysis interval and $n_{i}$ is the counts of $i^{\rm th}$ experimental data point. $p_{i}$ represents the probability that the theoretical model predictions falling within the corresponding histogram bin of the experimental data under the specified statistical analysis interval and for a given $R_{\text{HMT}}$. Additionally, in the given statistical interval, all the $p_{i}$ should be normalized. The second and third terms are model independent, which can be neglected to simplify the calculation. Thus, a simplified logarithmic likelihood function can be defined as
@@
To determine the confidence intervals for the HMT ratio based on the likelihood function distribution, $R_{\text{HMT}}$ values were identified that correspond to relative likelihood values smaller than the maximum by $\frac{m^{2}}{2}$, which represent confidence levels of $\pm m\sigma$. The $R_{\text{HMT}}$ ranges for likelihood values reduced by $0.5$, $4.5$ and $12.5$ from the quadratic maximum correspond to $\pm 1\sigma$, $\pm 3\sigma$ and $\pm 5\sigma$ confidence intervals, respectively, as indicated by the black dashed lines in the figure. Based on the experimentally measured $\gamma$ spectrum, the value of $R_{\text{HMT}}$ is determined at $1\sigma$ confidence level as
@@
V.3 Experimental $\gamma$ spectrum in fast coincidence
Alternatively, one can analyze the $\gamma$ spectrum in a different approach as a cross check, where both the $\gamma$ spectrum and the background are extracted in beam-on experiment. Specifically, the $\gamma$ events are accumulated in fast coincidence gate defined by the trigger, while the background is obtained in a time window with the same width but far away from the main trigger time window. By this approach, one ensures that only random coincidence are contained in the time window far away from the main trigger gate.
Noticeably, however, the events triggered by the SSD M2  condition shall be excluded now, because the $\gamma$ time are spreading across the entire TDC range due to serious timing jitter of the trigger signal. Otherwise, the truly coincident $\gamma$ events are recorded in the background window and the subtraction leads to a wrong energy spectrum.
To isolate true $\gamma$ events, we exclude all events of the SSD M2  trigger within the main peak region and define two time windows based on Fig. 13:

### v2

@@ III.1 IBUU-MDI Model Overview
For comparison, a momentum-independent soft Bertsch-Kruse-Das Gupta (SBKD) potential is also used in IBUU simulations [9],
@@
The bremsstrahlung reaction channel is incorporated into the collision processes within the IBUU model, which employs the Bertsch criterion [92] and Pauli exclusion principle to determine whether a NN collision occurs. Bremsstrahlung photons are generated through two channels, the neutron-proton collisions ($np\rightarrow np\gamma$) and the proton-proton collisions ($pp\rightarrow pp\gamma$). However, the $np\gamma$ channel dominates as the lowest possible multipole radiation in the $pp$ collisions is quadrupole, whose intensity is much lower compared to the dipole radiation in $np$ collisions [11, 54]. Hence, we consider only the $np\gamma$ process in our calculations.
The bremsstrahlung photons emitted in the $np$ collisions are produced in the processes of deceleration and acceleration of the protons interacting with the neutrons. In the intermediate energy range, we adopt the neutral scalar $\sigma$ meson exchange model [11] in the present work, in which photons are emitted via the external processes, illustrated by the Feynman diagrams in Fig.2 (a). At higher energies, one will have to take the charged meson exchange processes into consideration. The photons can also be emitted from the internal mesons, to which the electromagnetic field also couples [11, 68], and the Feynman diagram is given by Fig.2 (b).
@@
Due to the low probability of producing bremsstrahlung photons in the reaction, the impact of bremsstrahlung on nucleon kinematics is also negligible, and hence the perturbation method is applied to calculate the probability of photon production in each $np$ collision. Based on the $\sigma$ meson exchange model mentioned above, a well fitted expression for the probability of elementary double differential photon production can be applied in IBUU simulations [26],
@@
where $E_{\gamma}$ represents the energy of produced photons, and $E_{\rm max}$ represents the total available energy in the center-of-mass frame of the colliding neutron-proton pair. The coefficient $\alpha=0.7319-0.5898\beta$, where $\beta$ represents the velocity of the nucleon. The total photon production probability per event is the sum of the probabilities of all $np$ collisions producing photons throughout the entire process.
III.2 Sensitivity to Model Parameters
@@
In the IBUU simulations, one has to choose several important inputs which may influence the results of the final observables. Here we systematically check the sensitivity of bremsstrahlung photon production on the choices of parameters, namely the impact parameter, the nuclear mean-field potential, and the symmetry energy parameter.
@@
Figure 3: (a) The double differential photon production calculated by varying impact parameters. The impact parameters are taken to be $b=0\sim 5\,\rm{fm}$, $b=0\sim 9\,\rm{fm}$ and fixed values $b=0,5,9\,\rm{fm}$. The results using $b=0\sim 9,0,5,9\,\rm{fm}$ are normalized to the $b=0\sim 5\,\rm{fm}$ case. (b) The double differential photon production calculated by varying the choice of nuclear mean-field potentials. We compare the results from the MDI potential with those from the momentum-independent SBKD potential. The results from the SBKD potential are normalized to the MDI results. (c) The double differential photon production calculated by varying symmetry energy parameters. Three different $x$ parameters are employed in the simulations, namely x=1x=-1, $x=0$ and $x=1$. The results using $x=0$ and $x=1$ are normalized to the x=1x=-1 case.
Firstly, we show in Fig. 3(a) the variations of bremsstrahlung photon production by using different impact parameters. We adopt $b=0\sim 5\,\rm{fm}$ in our IBUU simulations. To check the sensitivity of photon production on $b$ parameter, several other values are used, namely, $b=0\sim 9\,\rm{fm}$ and $b=0,5,9\,\rm{fm}$. Their corresponding spectra are normalized to the $b=0\sim 5\,\rm{fm}$ one. This is because we focus on the spectral shape similarities rather than absolute normalization. Although the absolute values of bremsstrahlung photon production spectra vary with different choices of impact parameters, but the spectral shape remains almost the same to the impact parameters, as seen in Fig. 3(a).
In Fig. 3(b), the variations of bremsstrahlung photon production by using different nuclear mean-field potentials are shown. We adopt the MDI potential with the symmetry energy parameter x=1x=-1 [64]. The momentum-independent SBKD potential is used to check the sensitivity of mean-field potentials, and the parameters used in SBKD potential are listed in Table1. The $\sigma$ parameter represents the softness or stiffness of the potential with the variation of the nuclear density [9]. For the possible influence of symmetry energy parameters in MDI, we adopt three different parameters x=1x=-1, $x=0$ and $x=1$ in simulations (see Fig. 3(c)). The tests have shown that within the energy region investigated in this work, our conclusions are not affected by the choice of nuclear potentials or the symmetry energy parameters. This is probably because the bremsstrahlung photons interact with nucleons only electromagnetically, and are not sensitive to the isoscalar and isovector parts of nuclear interactions.
@@
Table 1: The parameters $\sigma$, AA (MeV) and $B$ (MeV) used in the momentum-independent SBKD potential and their corresponding incompressibility coefficients $K$ (MeV) in the equation of state of nuclear matter.
$\sigma$ AA (MeV) $B$ (MeV) $K$ (MeV)
7/6 -356 303 201
4/3 -218 164 237
2 -124 70.5 377
2.77 -103.22 49.56 540
IV Experimental $\gamma$-Ray Spectra
IV.1 Analysis of Systematic Uncertainty
@@
The systematic uncertainties primarily stem from three sources: (1) variations in the spectra obtained from radioactive source calibrations on different dates during the experiment, (2) the choice of normalization energy range for high-energy cosmic rays during background subtraction, and (3) the linear assumption of the detector response at higher $\gamma$ energy levels. Analysis shows that the choice of normalization energy ranges exhibits no significant variations, indicating that the effect of cosmic ray normalization range on the $\gamma$-ray spectrum is negligible.
@@
The experimental data included three sets of linear calibrations obtained from radioactive sources on different dates. Among these, the $\gamma$ energy spectrum calibrated using the dataset with the highest statistics was designated as the central spectrum. In addition to the linear calibration, we introduced two alternative detector response correction functions (DRCFs) characterizing the differences between the linear calibration and the actual quadratic calibration obtained on SLEGS test experiment [97]. Then a total of nine distinct $\gamma$ energy spectra can be derived. Excluding the central spectrum, we calculated the standard deviation of the remaining eight spectra for each bin as the systematic uncertainty. During the evaluation of the likelihood functions and $\chi^{2}$ (see next), the values derived from the central spectrum were used as the central points. The likelihood function and $\chi^{2}$ values computed from the remaining spectra were utilized to determine the associated errors (standard deviations). This methodology enabled us to obtain likelihood and $\chi^{2}$ functions with systematic uncertainties at different $R_{\text{HMT}}$ values.
IV.2 Experimental $\gamma$ spectrum in slow coincidence
@@
As shown in the primary analysis, the method introduced in the previous experiment [64] is applied. All $\gamma$ events in the slow coincidence window are counted. Namely a valid timing of the reconstructed $\gamma$ is found within the TDC range. The background is obtained by the beam-off measurement. The experimental data calibration and background subtraction were performed using the background data and radioactive source calibration data on March 8, 2024. Calibration files were generated based on these data and subsequently applied to all experimental and background data for calibration and reconstruction. This process yielded the total beam-on spectrum and the beam-of background spectrum, shown as the black and red lines in Fig. 4(a), respectively. The red background spectrum was scaled by a normalization factor $r_{\rm n}$. Specifically, the total counts above $110~\rm MeV$ in both the total spectrum ($Y_{\rm on}$) and the background spectrum ($Y_{\rm off}$) were calculated, and the yield ratio is defined as $r_{\rm n}=\frac{Y_{\rm on}(E_{1})}{Y_{\rm off}(E_{1})}$, where the subscript ‘on’ and ‘off’ represent the beam-on and beam-off spectrum, respectively. $E_{1}=110$ MeV is the low-energy border taken to do the normalization, and the upper border is set to 200 MeV. Then the ratio $r_{\rm n}$ is applied to the entire background spectrum, resulting in the red histogram in Fig. 4(a). By subtracting the scaled background spectrum from the total energy spectrum, the background-subtracted $\gamma$ spectrum was obtained, as shown in Fig. 4(b), where we take 01000-100 MeV as the energy range of the total $\gamma$-rays.
@@
Fig. 5 illustrates the rebinned experimental $\gamma$ spectrum in the c.m. frame. The black dots represent the central spectrum, with statistical uncertainties indicated by error bars, and the gray shaded regions denote the systematic uncertainties at each energy point. Several key theoretical curves, processed with the detector filter, are presented for comparison. Unlike in our previous measurement [64, 95], in this experiment, the $\gamma$ hodoscope was operated in active triggering mode, preventing the determination of the number of $np$ collisions in heavy-ion reactions as a normalization factor for the $\gamma$ spectrum. Consequently, the comparison between the experimental data and theoretical curves predicted by the IBUU model focused on spectral shape similarities rather than absolute normalization. To align the range of theoretical curves with the experimental spectrum, the theoretical values were uniformly scaled by a factor of $2.5\times 10^{9}$, which had no impact on determining the $R_{\rm HMT}$, as it did not affect the comparison of the shape differences between the experimental and model-predicted spectra.
Figure 5: (Color Online) Comparison of the rebinned experimental $\gamma$ spectrum (black dots) with statistical uncertainties (error bars) and systematic uncertainties (gray shaded areas) in the c.m. frame. Several key theoretical curves with different $R_{\rm HMT}$, processed with the detector filter, are overlaid for comparison.
Energies below $30~\rm MeV$ are primarily influenced by collective resonance and statistical emissions, as well as the random coincidence. Therefore, we selected an energy range of $35<E_{\gamma}^{\rm cm}<100~\rm MeV$ as the central analysis interval and calculated the likelihood function between the experimental $\gamma$ spectrum at each point within this range and various IBUU theoretical curves.
@@
where $i$ represents the sum of experimental points within a certain statistical analysis interval and $n_{i}$ is the count of $i^{\rm th}$ experimental data point. $p_{i}$ represents the probability that the theoretical model predictions falling within the corresponding histogram bin of the experimental data under the specified statistical analysis interval and for a given $R_{\text{HMT}}$. Additionally, in the given statistical interval, all the $p_{i}$ should be normalized. The second and third terms are model independent, which can be neglected to simplify the calculation. Thus, a simplified logarithmic likelihood function can be defined as
@@
To determine the confidence intervals for the HMT ratio based on the likelihood function distribution, $R_{\text{HMT}}$ values were identified that correspond to relative likelihood values smaller than the maximum by $\frac{m^{2}}{2}$, which represent confidence levels of $\pm m\sigma$. The $R_{\text{HMT}}$ ranges for likelihood values reduced by $0.5$, $4.5$ and $12.5$ from the quadratic maximum correspond to $\pm 1\sigma$, $\pm 3\sigma$ and $\pm 5\sigma$ confidence intervals, respectively, as indicated by the black dashed lines in the figure. Based on the experimentally measured $\gamma$ spectrum, the value of $R_{\text{HMT}}$ is determined at the $1\sigma$ confidence level as
@@
IV.3 Experimental $\gamma$ spectrum in fast coincidence
Alternatively, one can analyze the $\gamma$ spectrum in a different approach as a cross check, where both the $\gamma$ spectrum and the background are extracted in beam-on experiment. Specifically, the $\gamma$ events are accumulated in the fast coincidence gate defined by the trigger, while the background is obtained in a time window with the same width but far away from the main trigger time window. By this approach, one ensures that only random coincidence is contained in the time window far away from the main trigger gate.
Noticeably, however, the events triggered by the condition requiring two or more LCPs detected in the SSDTs (SSD M2) shall be excluded now, because the $\gamma$ time are spreading across the entire TDC range due to serious timing jitter of the trigger signal (see Supplementary Materials). Otherwise, the truly coincident $\gamma$ events are recorded in the background window and the subtraction leads to a wrong energy spectrum.
To isolate true $\gamma$ events, we exclude all events of the SSD M2  trigger within the main peak region and define two time windows: a main window and a window well separated from the former, as specified below,